# Bayesian Concept Bottleneck Models with LLM Priors

### v2

@@ C.3 Example prompt for annotations for candidate concepts (Step 3)
Note that in cases where the LLM was unsure about the concept’s value, it was allowed to return a probability, which prior works have found to be helpful [6]. In general, the LLM outputted 1’s and 0’s. Probabilities were outputted occasionally, which as highlighted in the main manuscript, can be audited and checked by a human.
@@ D.1 MIMIC
We evaluate BC-LLM against baseline methods using Cohere’s Command-R LLM using the same settings as discussed in Section 4.2. Performance can be found in Figure 5.
Figure 5: Comparison of BC-LLM and existing methods in terms of performance and recovery of true concepts. Error bars indicate the 95% CI.
@@ D.2 CUB-Birds dataset
We evaluate BC-LLM against baseline methods using Anthropic’s Claude 3.5 Haiku using the same settings as discussed in Section 4.1. Performance can be found in Table 2.
@@
Table 2: Results for the CUB-Birds dataset using Claude 3.5 Haiku
@@ E.3 CUB-Birds dataset
E.4 Clinical notes and tabular data from Hospital Anonymous
@@ Appendix F Laplace Approximation of Split-Sample Posterior
In the current implementation of our procedure, we use a slightly different method to perform Laplace approximation, which instead approximates the integral in (1).
@@
Step 2: Reduction to connected components. Since $(\tilde{\vec{c}}^{~t})_{t=1}^{\infty}$ satisfies the detailed balance equations, it is a symmetric Markov chain, i.e. it has an equivalent representation as a flow on a network, in which the vertices are states and the edges correspond to nonzero transition probabilities. This property is inherited by $(\vec{c}^{~t})_{t=1}^{\infty}$. Consider a connected component $\mathcal{A}$ of the network. The Markov chain, when restricted to $\mathcal{A}$, is aperiodic, since there must exist a self-loop. Furthermore, it is finite according to Assumption G.1(vi). As such, $(\vec{c}^{~t})_{t=1}^{\infty}$ has a unique stationary distribution $\pi_{\mathcal{A}}$ supported on $\mathcal{A}$ [55]. Indeed, all stationary distributions of $(\vec{c}^{~t})_{t=1}^{\infty}$ are convex combinations of such distributions. Hence, it suffices to show the desired property for $\pi_{\mathcal{A}}$.
@@
We can thus use Talagrand’s comparison inequality [56] to get, for any $t>0$, that
@@ Appendix H Stationary Distribution Under Consistent Proposals
as we wanted. ∎

### v3

@@ C.3 Example prompt for annotations for candidate concepts (Step 3)
Note that in cases where the LLM was unsure about the concept’s value, it was allowed to return a probability, which prior works have found to be helpful [Kim2023-vu]. In general, the LLM outputted 1’s and 0’s. Probabilities were outputted occasionally, which as highlighted in the main manuscript, can be audited and checked by a human.
@@ D.1 MIMIC
To assess the robustness of BC-LLM, we include the following variations. First, we test the robustness of BC-LLM to prompt phrasing. In particular, we feed in a prompt that is more vague than that used in the main manuscript during Step 0 of BC-LLM by replacing the original sentence “Output a list of descriptors that summarizes the patient case (such as aspects on demographics, diagnoses, SDOH, etc).“ with “Output a list of descriptors that summarizes the patient case.“ Second, we test the robustness of the results to the choice of LLM by rerunning the same experiment using Cohere’s Command-R LLM, rather than GPT-4o-mini.
Results are shown in Figure 5. First, we see that BC-LLM is quite robust to prompt phrasing, as the results with this simpler prompt are similar to the more detailed prompt. Second, we see that performance of all the methods—BC-LLM and the comparator methods—are lower using Command-R. Nevertheless, we see that the relative performance between the methods is the same, with BC-LLM performing the best. [UPDATE FIGURE – BC-LLM should also be labelled with cohere vs prompt ablation]
Figure 5: Additional results running BC-LLM and comparator methods on the MIMIC dataset, evaluated in terms of performance and recovery of true concepts. Error bars indicate the 95% CI.
@@ D.2 CUB-Birds dataset
Table 2 reports results from two additional sets of experiments on the CUB-Birds dataset. First, to evaluate robustness of BC-LLM to the choice of LLM, we report the performance of BC-LLM and comparator methods when Anthropic’s Claude 3.5 Haiku is used instead, with all other settings the same as that discussed in Section 4.1. Second, to facilitate comparisons to other CBM results on the CUB-Birds dataset, we evaluate BC-LLM on the dataset’s original task of classifying all 200 bird species. Nevertheless, we emphasize that this is not the intended use case of BC-LLM, as its aim is to learn fully interpretable CBM methods with $K\leq 20$. Due to computation time, we run BC-LLM for one epoch rather than five. [technically it is half an epoch…]
@@
Classifying between all 200 bird species
BC-LLM 0.562 (0.487, 0.644) 0.598 (0.523, 0.665)
LLM+CBM
Boosting LLM+CBM 0.005 (0.003, 0.007) 0.500 (0.500, 0.500) 1.012 (1.011, 1.012)
Table 2: Additional results for the CUB-Birds dataset
D.3 Other image datasets
To evaluate BC-LLM on other real-world image datasets, Table 3 includes two additional imaging datasets:
  • Functional Map of the World (fMoW) [Christie2018-dm]: RGB satellite images classified into 62 building/land use categories. Models were trained on satellite images from the US, with 100 images sampled from each category. CBMs were trained to have 60 concepts. Models were evaluated in terms of in-distribution performance (i.e. images from the USA) as well as out-of-distribution (OOD) performance when evaluated on satellite images from China.
  • Imagenette [imagenette-xz]: 10 classes from ImageNet. CBMs were trained to have 10 concepts.
Method Accuracy ($\uparrow$) AUC ($\uparrow$) Brier ($\downarrow$)
fMoW: USA images
BC-LLM 0.357 (0.340, 0.375) 0.904 (0.898, 0.910) 0.780 (0.766, 0.792)
LLM+CBM 0.311 (0.295, 0.327) 0.878 (0.871, 0.885) 0.892 (0.872, 0.914)
Boosting LLM+CBM 0.118 (0.105, 0.130) 0.757 (0.748, 0.764) 1.029 (1.018, 1.039)
fMoW: OOD-China images
BC-LLM 0.265 (0.246, 0.285) [???] 0.853 (0.839, 0.868)
LLM+CBM 0.223 (0.205, 0.242) [???] 1.023 (1.00, 1.045)
Boosting LLM+CBM 0.084 (0.072, 0.098) [??] 1.065 (1.054, 1.077)
Imagenette
BC-LLM 0.987 (0.980, 0.993) 0.999 (0.997, 1.000) 0.072 (0.062, 0.085)
LLM+CBM 0.902 (0.883, 0.919) 0.988 (0.986, 0.990) 0.125 (0.107, 0.142)
Boosting LLM+CBM 0.639 (0.610, 0.670) 0.899 (0.888, 0.909) 0.448 (0.422, 0.479)
Table 3: Results on additional imaging datasets
@@ E.3 CUB-Birds dataset
E.4 Clinical notes and tabular data from the Zuckerberg San Francisco General Hospital
@@ Appendix F Laplace Approximation of Split-Sample Posterior
In the current implementation of our procedure, we use a slightly different method to perform Laplace approximation, which instead approximates the integral in (LABEL:eq:conditional_lkhd_decomposition).
@@
Step 2: Reduction to connected components. Since $(\tilde{\vec{c}}^{~t})_{t=1}^{\infty}$ satisfies the detailed balance equations, it is a symmetric Markov chain, i.e. it has an equivalent representation as a flow on a network, in which the vertices are states and the edges correspond to nonzero transition probabilities. This property is inherited by $(\vec{c}^{~t})_{t=1}^{\infty}$. Consider a connected component $\mathcal{A}$ of the network. The Markov chain, when restricted to $\mathcal{A}$, is aperiodic, since there must exist a self-loop. Furthermore, it is finite according to Assumption G.1(vi). As such, $(\vec{c}^{~t})_{t=1}^{\infty}$ has a unique stationary distribution $\pi_{\mathcal{A}}$ supported on $\mathcal{A}$ [LevinPeresWilmer2006]. Indeed, all stationary distributions of $(\vec{c}^{~t})_{t=1}^{\infty}$ are convex combinations of such distributions. Hence, it suffices to show the desired property for $\pi_{\mathcal{A}}$.
@@
We can thus use Talagrand’s comparison inequality [vershynin2018high] to get, for any $t>0$, that
@@ Appendix H Stationary Distribution Under Consistent Proposals
as we wanted.
To show the statement for Multi-SS-MH-Update, we use its equivalence, for a fixed $S$, to the modified Multiple-Try Metropolis-Hastings method described in Appendix A. In Appendix A, we prove that this method satisfies the detailed balance equations for the posterior. Since the detailed balance equations are linear in the transition probabilities, we can marginalize this over $S$ to show that the posterior is stationary with respect for Multi-SS-MH-Update, which draws $S$ uniformly at random. ∎